Ronald was recently hired by Highland Equipment Inc. as a junior budget analyst. He is working for the Venture Capital Division and has been given for capital budgeting projects to evaluate. He must give his analysis and recommendation to the capital budgeting committee.

Ronald has a B.S. in accounting from (2011) and passed the CPA exam (2017). He has been in public accounting for several years. During that time he earned an MBA from Seattle U. He would like to be the CFO of a company someday--maybe Highland Equipment Inc.-- and this is an opportunity to get onto that career track and to prove his ability.

As Ronald looks over the financial data collected, he is trying to make sense of it all. He already has the most difficult part of the analysis complete -- the estimation of cash flows. Through some internet research and application of finance theory, he has also determined the firm’s beta.

Here is the information that Ronald has accumulated so far:

The Capital Budgeting Projects

He must choose one of the four capital budgeting projects listed below:  

Table 1

t

A

B

C

D

0

      (19,000,000)

      (20,000,000)

      (14,000,000)

       (18,000,000)

1

         5,200,000

         6,000,000

         5,200,000

          7,600,000

2

         5,200,000

       10,000,000

         5,200,000

          7,600,000

3

         6,800,000

         8,000,000

         5,200,000

          5,600,000

4

         6,800,000

         4,000,000

         5,200,000

          5,600,000

Risk

Average

High

Low

Average

Table 1 shows the expected after-tax operating cash flows for each project. All projects are expected to have a 4 year life. The projects differ in size (the cost of the initial investment), and their cash flow patterns are different. They also differ in risk as indicated in the above table.

The capital budget is $20 million and the projects are mutually exclusive.

Capital Structures

Highland Equipment Inc. has the following capital structure, which is considered to be optimal:

Debt  

40%

Preferred Equity

5%

Common Equity

55%

100%

   

Cost of Capital

Ronald knows that in order to evaluate the projects he will have to determine the cost of capital for each of them. He has been given the following data, which he believes will be relevant to his task.

(1)The firm’s tax rate is 35%.

(2) Highland Equipment Inc. has issued a 12% semi-annual coupon bond with 5 years term to maturity. The current trading price is $1,040.

(3) The firm has issued some preferred stock which pays an annual 11% dividend of $100 par value, and the current market price is $106.

(4) The firm’s stock is currently selling for $95 per share. Its last dividend (D0) was $5, and dividends are expected to grow at a constant rate of 7%. The current risk free return offered by Treasury security is 1.5%, and the market portfolio’s return is 9.5%. Highland Equipment Inc. has a beta of 1.4. For the bond-yield-plus-risk-premium approach, the firm uses a risk premium of 3%.

(5) The firm adjusts its project WACC for risk by adding 2% to the overall WACC for high-risk projects and subtracting 2% for low-risk projects.

Ronald knows that Highland Equipment Inc. executives have favored IRR in the past for making their capital budgeting decisions. His professor at Seattle U. said NPV was better than IRR. His textbook says that MIRR is also better than IRR. He is the new kid on the block and must be prepared to defend his recommendations.

First, however, Ronald must finish the analysis and write his report. To help begin, he has formulated the following questions:

  1. What is the firm’s cost of debt?
  1. What is the cost of preferred stock for Highland Equipment Inc.?
  1. Cost of common equity

(1) What is the estimated cost of common equity using the CAPM approach?

(2) What is the estimated cost of common equity using the DCF approach?

(3) What is the estimated cost of common equity using the bond-yield-plus-risk-premium approach?

(4) What is the final estimate for rs?

  1. What is Highland Equipment Inc.’s overall WACC?
  1. Do you think the firm should use the single overall WACC as the hurdle rate for each of its projects? Explain.
  1. What is the WACC for each project? Place your numerical solutions in Table 2.
  1. Calculate all relevant capital budgeting measures for each project, and place your numerical solutions in Table 2.

Table 2

A

B

C

D

WACC

NPV

IRR

MIRR

  1. Comment on the commonly used capital budgeting measures. What is the underlying cause of ranking conflicts? Which criterion is the best one, and why?
  1. Which of the projects are unacceptable and why?
  1. Rank the projects that are acceptable, according to Ronald’s criterion of choice.
  1. Which project should Ronald recommend and why? Explain why each of the projects not chosen was rejected.
Answer #1

What is the firm’s cost of debt?
Pre-tax cost=RATE(5*2,12%*1000/2,-1040,1000)*2=10.94%

After-tax cost=RATE(5*2,12%*1000/2,-1040,1000)*2*(1-35%)=7.111%

What is the cost of preferred stock for Highland Equipment Inc.?
=11%*100/106=10.377%

Cost of common equity
(1) What is the estimated cost of common equity using the CAPM approach?
=1.5%+1.4*(9.5%-1.5%)=12.700%

(2) What is the estimated cost of common equity using the DCF approach?
=5*(1+7%)/95+7%=12.632%

(3) What is the estimated cost of common equity using the bond-yield-plus-risk-premium approach?
=RATE(5*2,12%*1000/2,-1040,1000)*2+3%=13.940%

(4) What is the final estimate for rs?
=(12.70%+12.632%+13.940%)/3=13.091%

P.S.: I am not allowed to answer more than 4 questions